OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (1, –7); Slope:-2/3 A. y –7 = (x – 1) B. y + 7 = (x + 1) C. y + 7 = (x – 1) D. y –7 = (x + 1

OpenStudy (♪chibiterasu):

Point: (1, –7); Slope:-2/3 y = mx + b m is the slope I assume the point is on the y -axis?

OpenStudy (♪chibiterasu):

If so, it is plugged into b.

OpenStudy (♪chibiterasu):

Waaiit point-slope not slope-intercept

OpenStudy (♪chibiterasu):

point-slope is: (y - y) = m(x - x) I think . . .

OpenStudy (♪chibiterasu):

So: y + 7 = -2/3(x - 1)

OpenStudy (♪chibiterasu):

So the closest answer is A, since none of them include the slope.

OpenStudy (♪chibiterasu):

Unless the slope is in there it just didn't appear when you pasted it. So yeah, A.

OpenStudy (anonymous):

The slope is....

OpenStudy (♪chibiterasu):

m, which is -2/3

OpenStudy (♪chibiterasu):

It's A lol

OpenStudy (anonymous):

This is the slope

OpenStudy (♪chibiterasu):

I know that lol xD It's A

OpenStudy (anonymous):

Alright, thanks.

OpenStudy (♪chibiterasu):

pff, ok x3